# The Influence of Exclusion Zones on the Coexistence of Predator and Prey with an Allee Effect

### v1

@@ 3. Coexistence equilibrium in higher dimensions
$\begin{cases}-d_{u}\Delta(w-1)+K(w-1)\leq 0\quad\text{in}\quad\Omega,\\ \partial_{\nu}(u-1)=0\quad\text{on}\quad\partial\Omega,\end{cases}$
whence, by the maximum principle, $u\leq 1$ on $\overline{\Omega}$. From the mast inequality we infer that $u\not\equiv 1$ and thus $u<1$ on $\overline{\Omega}$. Likewise, since $d_{u}\Delta w+Kw\geq 0$ on $\overline{\Omega}$ and $d_{u}\Delta(w-\zeta)+K(w-\zeta)\geq 0$ on $\Omega\setminus A$, we get $w\geq 0$ on $\overline{\Omega}$ and $w\geq\zeta$ on $\overline{\Omega}\setminus B_{\rho}$. From the fact that $\zeta$ extended by 0 on $B_{\rho}$ is a strict sub-solution of the equation, we derive that these inequalities are strict and therefore that $w$ is in the interior of $D$. ∎
@@
which establishes the claim of the proposition 3.10. ∎
@@
4. Persistence owing to exclusion zones
Theorem 1.3 shows that the 1D stationary model (2) has solutions provided LaL-a is large enough. Extending this, Theorem 1.4 shows that the higher dimensional model (1) has coexistence equilibria provided $\Omega\setminus A$ contains a large enough ball. In other words, if the prey have a large predator-free exclusion zone, then both species can persist. In this section we show that this is a necessary condition.
@@ 4. Persistence owing to exclusion zones
if and only if $\theta<\gamma/\alpha<1$ because otherwise, either $\hat{u}\not\in(0,1)$ or f(u^)f($\hat${u}) (and hence $\hat{v}$) is negative. This suggests that for small $\gamma/\alpha$, the equilibrium solutions of (1) may fail to be coexistenece solutions due since, for such solutions $(u,v)$, we must have $f(u)<0$. In other words, the overabundance of predators resulting from their high predation efficiency $\alpha$ and low mortality rate $\alpha$ forces the prey population beneath the Allee threshold. This fact is proved by the following theorem.
@@
Suppose $A=\Omega$, a bounded domain in $\mathbb{R}^{N}$ with $C^{2}$ boundary. There exists $\gamma_{0}>0$ (depending on $\alpha,\beta,f$, and $\Omega$) such then if $0<\gamma<\gamma_{0}$, then (1) does not have a coexistence equilibrium.
@@
Fix $\alpha$, $\beta$, $f$, and $\Omega\subset\mathbb{R}^{N}$ with $\Omega$ bounded. Suppose by way of contradiction that there exists a coexistence equilibrium $(u_{n},v_{n})$ to (1) with $A=\Omega$ and $\gamma_{n}=1/n$ for each $n\in\mathbb{N}$. As a coexistence equilibrium, $(u_{n},v_{n})$ satisfies $0<u_{n}<1$ and $0<v_{n}$.
@@
$\beta\left(\inf_{\Omega}v_{n}\right)\int_{\Omega}u_{n}^{2}\leq D_{u}\|\nabla u_{n}\|_{L^{2}}^{2}+\beta\int_{\Omega}v_{n}u_{n}^{2}=\int_{\Omega}u_{n}f(u_{n})\leq L\int_{\Omega}u_{n}^{2},$ (29)
where $L=\sup_{0<s\leq 1}\tfrac{f(s)}{s}$. Therefore, for all $n$, $\inf_{\Omega}v_{n}\leq\frac{L}{\beta}$.
On the other hand, $v_{n}$ satisfies the elliptic equation $D_{v}\Delta v_{n}+\left(\alpha u_{n}-\frac{1}{n}\right)v_{n}=0$. By Harnack’s Inequality, there exists a constant $C_{n}>0$ depending on $\Omega$ and $\|\alpha u_{n}-1/n\|_{L^{\infty}}$ such that $\sup_{\Omega}v_{n}\leq C_{n}\inf_{\Omega}v_{n}$. But since $\|\alpha u_{n}-1/n\|_{L^{\infty}}$ is bounded, so is $C_{n}$, so there is a single constant $C>0$ such that
$\|v_{n}\|_{L^{\infty}}=\sup_{\Omega}v_{n}\leq C\inf_{\Omega}v_{n}.$ (30)
We conclude that for all $n$, $\|v_{n}\|_{L^{\infty}}\leq\frac{LC}{\beta}$. Thus, $\|v_{n}\|_{L^{\infty}}$ is bounded.
@@
then if $g\in L^{p}(\Omega)$, then $w$ is in $W^{2,p}(\Omega)$ and there is a constant $K$ (depending only on $\Omega$) such that $\|w\|_{W^{2,p}}\leq K\|g\|_{L^{p}}$ (see, e.g., Theorem 9.26 in [9]). For each $n$, we may write the equations for $u_{n}$ and $v_{n}$ as
@@
Since Harnack’s inequality (30) applies to $\tilde{v}_{n_{k}}$ for each $k$, it also applies to $\tilde{v}_{0}$. But we have $\sup_{\Omega}\tilde{v}_{0}=1$, so $\tilde{v}_{0}>0$ in $\Omega$. Thus, $u_{0}\equiv 0$. But $u_{n_{k}}$ converges to $u_{0}$ uniformly, meaning that for $k$ sufficiently large, $0<u_{n_{k}}<\theta$ in $\Omega$. But for such $k$,
@@
It is worth noting how this proof breaks down if $A$ is a proper subset of $\Omega$. In the latter case, we may still find limits $u_{0}$ and $v_{0}$ such that $\int_{A}u_{0}v_{0}=0$, and conclude that $u_{0}=0$ on $A$, but we may not conclude that $u_{0}=0$ on $\Omega$. This is why coexistence equilibria may still be guaranteed by Theorem 1.4.
5. Thin-limit analysis in 1D
In this section, we examine the behavior of solutions to (1) in the limit as the predator domain becomes small. Specifically, we will consider $\Omega=(0,L)$ and $A=(0,a)$ as $a\to 0^{+}$.
One may expect that a predator domain of size zero cannot support any predators, but this is not necessarily the case. Consider, for example, crocodiles hunting for prey along the shore of a watering hole. Indeed, our first result on this subject below is that as $a\to 0$, the total predator population approaches a positive limit.
@@ 5. Thin-limit analysis in 1D
Suppose $\alpha>\gamma$. Fix $a_{0}>0$ and let $L_{0}$ be sufficiently large so that for all $L>L_{0}$ and $a<a_{0}$, Theorem 1.3 holds. For fixed $L>L_{0}$ and for each $0<a<a_{0}$, let $u_{a}$ and $v_{a}$ denote a coexistence equilibrium to (1) with $\Omega=(0,L)$ and $A=(0,a)$ guaranteed by (1.3). Then
$0<\liminf_{a\to 0^{+}}\int_{0}^{a}v_{a}(x)\,dx<\limsup_{a\to 0^{+}}\lim_{a\to 0^{+}}v_{a}(x)\,dx<\infty.$ (36)
@@
By Theorem 1.3, for all $0<a<L-b_{0}$, $0<u_{a}<1$ and $0<v_{a}$. Let
$V_{-}=\liminf_{a\to 0+}\int_{0}^{a}v_{a}(x)\,dx\quad\text{and}\quad V_{+}=\limsup_{a\to 0+}\int_{0}^{a}v_{a}(x)\,dx.$ (37)
Clearly, $0\leq V_{-}\leq V_{+}\leq\infty$. We will show that in fact, $0<V_{-}$ and $V_{+}<\infty$.
Step 1. First we show that $0<V_{-}$. Suppose, by way of contradiction, that $V_{-}=0$. Then there exists a sequence $\{c_{n}\}\subset(0,L-b_{0})$ converging to 0 so that
$\lim_{n\to\infty}\int_{0}^{c_{n}}v_{c_{n}}(x)\,dx=0.$
For each $n$, $u_{c_{n}}$ and $v_{c_{n}}$ satisfy $v_{c_{n}}^{\prime\prime}+(\alpha u_{c_{n}}-\gamma)v_{c_{n}}$ in $(0,c_{n})$ with Neumann boundary conditions. Therefore, by the Poincaré inequality, there exists $C>0$ so that
$\begin{split}\|v_{c_{n}}^{\prime}\|_{L^{1}}&\leq C\|v_{c_{n}}^{\prime\prime}\|_{L^{1}}\\ &=C\|(\alpha u_{c_{n}}-\gamma)v_{c_{n}}\|_{L^{1}}\\ &\leq C\|\alpha u_{c_{n}}-\gamma\|_{L^{\infty}}\|v_{c_{n}}\|_{L^{1}}\\ &\leq C(\alpha+\gamma)\|v_{c_{n}}\|_{L^{1}}.\end{split}$ (38)
On the other hand, for any $x\in(0,c_{n})$
$|v_{c_{n}}(x)-v_{c_{n}}(0)|=\left|\int_{0}^{x}v_{c_{n}}^{\prime}(s)\,ds\right|\leq\int_{0}^{c_{n}}|v_{c_{n}}^{\prime}(s)|\,ds=\|v_{c_{n}}^{\prime}\|_{L^{1}}.$ (39)
Therefore, $\lim_{n\to\infty}\|v_{c_{n}}\|_{L^{\infty}}=0$. Furthermore, we have
$0=\int_{0}^{L}\left(u_{c_{n}}^{\prime\prime}+f(u_{c_{n}})-\beta\mathbb{1}_{[0,c_{n}]}u_{c_{n}}v_{c_{n}}\right)\,dx=\int_{0}^{L}f(u_{c_{n}})\,dx-\beta\int_{0}^{c_{n}}u_{c_{n}}v_{c_{n}}\,dx.$ (40)
We conclude that $\lim_{n\to\infty}\int_{0}^{L}f(u_{c_{n}})\,dx=0$. We may repeat this calculation while first multiplying by $u_{c_{n}}$ to obtain
$\begin{split}0&=\int_{0}^{L}\left(u_{c_{n}}^{\prime\prime}+f(u_{c_{n}})-\beta\mathbb{1}_{[0,c_{n}]}u_{c_{n}}v_{c_{n}}\right)u_{c_{n}}\,dx\\ &=-\|u_{c_{n}}^{\prime}\|_{L^{2}}^{2}+\int_{0}^{L}f(u_{c_{n}})u_{c_{n}}\,dx-\beta\int_{0}^{c_{n}}u_{c_{n}}^{2}v_{c_{n}}\,dx\end{split}$ (41)
Since $0<u_{n}<1$, we have $\lim_{n\to\infty}\int_{0}^{L}f(u_{c_{n}})u_{c_{n}}\,dx=0$ and $\lim_{n\to\infty}\int_{0}^{c_{n}}u_{c_{n}}^{2}v_{c_{n}}\,dx=0$. Therefore, $\lim_{n\to\infty}\|u_{c_{n}}^{\prime}\|_{L^{2}}=0$. This means that for any $x\in(0,L)$,
$|u_{c_{n}}(x)-u_{c_{n}}(0)|\leq\|u_{c_{n}}^{\prime}\|_{L^{1}}\leq L\|u_{c_{n}}^{\prime}\|_{L^{2}}.$ (42)
Thus, for large $n$, $u_{c_{n}}$ is close to $u_{c_{n}}(0)$. Since $0\leq u_{c_{n}}(0)\leq 1$, we may pass to a subsequence such that $u_{c_{n}}(0)$ converges to some $u_{0}\in[0,1]$. Therefore, $u_{c_{n}}$ converges to $u_{0}$ in $L^{\infty}$. Clearly, $f(u_{0})=0$, but since $\zeta(x)\leq u_{c_{n}}\leq 1$ for all $n$, the only candidate root of $f$ is $u_{0}=1$.
This means that for $n$ sufficiently large, $\alpha u_{c_{n}}-\gamma>0$. For such $n$, we have
$0=\int_{0}^{c_{n}}\left(v_{c_{n}}^{\prime\prime}+(\alpha u_{c_{n}}-\gamma)v_{c_{n}}\right)\,dx=\int_{0}^{c_{n}}(\alpha u_{c_{n}}-\gamma)v_{c_{n}}\,dx.$ (43)
But this is a contradiction because the integrand is positive. Therefore, we conclude that $V_{-}>0$.
Step 2. Now we show that $V_{+}<\infty$. Once again, there is a sequence $\{d_{n}\}\subset(0,L-b_{0})$ converging to 0 such that
$\lim_{n\to\infty}\int_{0}^{d_{n}}v_{d_{n}}\,dx=\infty.$ (44)
Then
$0=\int_{0}^{d_{n}}\left(v_{d_{n}}^{\prime\prime}+(\alpha u_{d_{n}}-\gamma)v_{d_{n}}\right)=\alpha\int_{0}^{d_{n}}u_{d_{n}}v_{d_{n}}\,dx-\gamma\int_{0}^{d_{n}}v_{d_{n}}\,dx.$ (45)
We conclude that $\lim_{n\to\infty}\int_{0}^{d_{n}}u_{d_{n}}v_{d_{n}}\,dx=\infty$. This means that
$\begin{split}0&=\int_{0}^{d_{n}}\left(u_{d_{n}}^{\prime\prime}+f(u_{d_{n}})-\beta u_{d_{n}}v_{d_{n}}\right)\,dx\\ &=u_{d_{n}}^{\prime}(d_{n})+\int_{0}^{d_{n}}f(u_{d_{n}})\,dx-\beta\int_{0}^{d_{n}}u_{d_{n}}v_{d_{n}}\,dx.\end{split}$ (46)
Since $f(u_{d_{n}})$ is uniformly bounded, we conclude that $\lim_{n\to\infty}u_{d_{n}}^{\prime}(d_{n})=\infty$. Finally,
$u_{d_{n}}^{\prime}(d_{n})=-\int_{d_{n}}^{L}u^{\prime\prime}_{d_{n}}\,dx=\int_{d_{n}}^{L}f(u_{d_{n}})\,dx.$ (47)
But, once again, $f$ is bounded on $[0,1]$, so $\int_{d_{n}}^{L}f(u_{d_{n}})\,dx$ is bounded, a contradiction. Therefore, $V_{+}<\infty$.
From Theorem 5.1, we see that as $a\to 0^{+}$, the total predator population of coexistence equilibria for (1) with $\Omega=(0,L)$ and $A=(0,a)$ neither vanishes nor blows up. The numerics in Section 6 suggest that, while there may not be a unique coexistence equilibrium for (1) in this setting, there is a unique asymptotically stable coexistence equilibrium. Moreover, if $v_{a}$ represents the predator component of this equilibrium, then the limit $\lim_{a\to 0^{+}}V(a)$ exists where $V(a)=\int_{0}^{a}v_{a}(x)\,dx$. Assuming that the $v_{a}$ and $V(a)$ indeed exist, we can compute asymptotics for $V(a)$ using the following theorem.
Theorem 5.2.
Suppose that the assumptions of Theorem 5.1 hold, and that in addition, the map $a\mapsto(u_{a},v_{a})$ is twice continuously differentiable with respect to the $C^{2}$ norm for $0<a<L-b_{0}$. Then
where $w_{1}$ and $w_{2}$ are solutions to
$\begin{cases}\frac{d_{u}}{L^{2}}w_{1}^{\prime\prime}+f(w_{1})=0&0<x<L\\ w_{2}^{\prime\prime}+w_{2}f^{\prime}(w_{1})=\frac{2}{L}f(w_{1})&0<x<L\\ w_{1}^{\prime}(L)=0&\\ w_{1}(0)=\frac{\gamma}{\alpha}&\\ w_{2}^{\prime}(L)=0&\\ w_{2}(0)=\frac{1}{3}w_{1}^{\prime}(0).\end{cases}$ (49)
@@
For any $a\in(0,L-b_{0})$, let $u_{1}=u_{a}\big|_{(0,a)}$ and $u_{2}=u_{a}\big|_{(a,L)}$. We apply the coordinate transformation $x\to x/a$ to $u_{1}$ and $v$ and the coordinate transformation $x\to(x-a)/(L-a)$ to $u_{2}$ to obtain equations for $u_{1}$, $u_{2}$, and $v$ in these new coordinates:

### v2

@@ 2 Coexistence equilibrium in higher dimensions
$\begin{cases}-d_{u}\Delta(w-1)+K(w-1)\leq 0\quad\text{in}\quad\Omega,\\ \partial_{\nu}(w-1)=0\quad\text{on}\quad\partial\Omega,\end{cases}$
whence, by the maximum principle, $u\leq 1$ on $\overline{\Omega}$. From the last inequality we infer that $u\not\equiv 1$ and thus $u<1$ on $\overline{\Omega}$. Likewise, since $d_{u}\Delta w+Kw\geq 0$ on $\overline{\Omega}$ and $d_{u}\Delta(w-\zeta)+K(w-\zeta)\geq 0$ on $\Omega\setminus A$, we get $w\geq 0$ on $\overline{\Omega}$ and $w\geq\zeta$ on $\overline{\Omega}\setminus B_{\rho}$. From the fact that $\zeta$ extended by 0 on $B_{\rho}$ is a strict sub-solution of the equation, we derive that these inequalities are strict and therefore that $w$ is in the interior of $D$. ∎
@@
which establishes the claim of Proposition 2.10. ∎
@@
3 Persistence owing to exclusion zones
Theorem 4.2 shows that the 1D stationary model (26) has solutions provided LaL-a is large enough. Extending this, Theorem 1.2 shows that the higher dimensional model (1) has coexistence equilibria provided $\Omega\setminus A$ contains a large enough ball. In other words, if the prey have a large predator-free exclusion zone, then both species can persist. In this section we show that the presence of an exclusion zone is, in some circumstances, necessary for coexistence solutions to exist..
@@ 3 Persistence owing to exclusion zones
if and only if $\theta<\gamma/\alpha<1$ because otherwise, either $\hat{u}\not\in(0,1)$ or f(u^)f($\hat${u}) (and hence $\hat{v}$) is negative. This suggests that for small $\gamma/\alpha$, the equilibrium solutions of (1) may fail to be coexistence solutions since, for such solutions $(u,v)$, we must have $f(u)<0$. In other words, the overabundance of predators resulting from their high conversion rate $\alpha$ and low mortality rate $\gamma$ forces the prey population beneath the Allee threshold. This fact is proved by the following theorem.
@@
Suppose $A=\Omega$, a bounded domain in $\mathbb{R}^{N}$ with $C^{2}$ boundary. There exists $\gamma_{0}>0$ (depending on $\alpha,\beta,f$, and $\Omega$) such that if $0<\gamma<\gamma_{0}$, then (1) does not have a coexistence equilibrium.
@@
Fix $\alpha$, $\beta$, $f$, and $\Omega\subset\mathbb{R}^{N}$ with $\Omega$ bounded. Suppose, by way of contradiction, that there exists a coexistence equilibrium $(u_{n},v_{n})$ to (1) with $A=\Omega$ and $\gamma_{n}=1/n$ for each $n\in\mathbb{N}$. Since $(u_{n},v_{n})$ is a coexistence equilibrium, it satisfies $0<u_{n}<1$ and $0<v_{n}$.
@@
$\beta\left(\inf_{\Omega}v_{n}\right)\int_{\Omega}u_{n}^{2}\leq d_{u}\|\nabla u_{n}\|_{L^{2}}^{2}+\beta\int_{\Omega}v_{n}u_{n}^{2}=\int_{\Omega}u_{n}f(u_{n})\leq Q\int_{\Omega}u_{n}^{2},$ (20)
where $Q=\sup_{0<s\leq 1}\tfrac{f(s)}{s}$. Therefore, for all $n$, $\inf_{\Omega}v_{n}\leq\frac{Q}{\beta}$.
On the other hand, by Lemma 2.8, there exists $\eta>0$ so that, regardless of $u_{n}$, a solution $\phi_{n}$ to
$\Delta\phi_{n}-(\alpha u_{n}-\gamma)\phi_{n}=0,\quad\quad\sup_{\Omega}\phi_{n}=1$
satisfies $0<\eta\leq\phi_{n}(x)\leq 1$ for all $x\in\Omega$. But $\phi_{n}=\frac{v_{n}}{\|v_{n}\|_{L^{\infty}}}$ is such a solution, so
$\eta\|v_{n}\|_{L^{\infty}}\leq\inf_{\Omega}v_{n}.$
We conclude that for all $n$, $\|v_{n}\|_{L^{\infty}}\leq\frac{Q}{\eta\beta}$. Thus, $\|v_{n}\|_{L^{\infty}}$ is bounded.
@@
and if $g\in L^{p}(\Omega)$, then $w$ is in $W^{2,p}(\Omega)$ and there is a constant $K$ (depending only on $\Omega$) such that $\|w\|_{W^{2,p}}\leq K\|g\|_{L^{p}}$ (see, e.g., Theorem 9.26 in [8]). For each $n$, we may write the equations for $u_{n}$ and $v_{n}$ as
@@
Since Lemma (2.8) applies to $\tilde{v}_{n_{k}}$ for each $k$, it also applies to $\tilde{v}_{0}$. But we have $\sup_{\Omega}\tilde{v}_{0}=1$, so $\tilde{v}_{0}>0$ in $\Omega$. Thus, $u_{0}\equiv 0$. But $u_{n_{k}}$ converges to $u_{0}$ uniformly, meaning that for $k$ sufficiently large, $0<u_{n_{k}}<\theta$ in $\Omega$. But for such $k$,
@@
It is worth noting how this proof breaks down if AA is a proper subset of $\Omega$. In the latter case, we may still find limits $u_{0}$ and $v_{0}$ such that $\int_{A}u_{0}v_{0}=0$, and conclude that $u_{0}=0$ on AA, but we may not conclude that $u_{0}=0$ on $\Omega$. This is why coexistence equilibria may still be possible via Theorem 1.2.
4 Thin-limit analysis in 1D
In this section, we consider the 1D steady state problem in the case that $\Omega=(0,L)$ and $A=(0,a)$ for $0<a<L$. That is, we study solutions to
[…]
with $v\not\equiv 0$.
We note that Theorem 1.2 does not apply to the setting (26) because AA and $\Omega$ share a common boundary at $x=0$. However, Theorem 4.2 below shows that the result of Theorem 1.2 may be extended to (26).
As in the higher-dimensional case, we will employ as a lower bound for $u$ in (26) the function $\zeta_{a,L}$ defined on $[a,L]$:
$\begin{cases}-d_{u}\zeta_{a,L}^{\prime\prime}=f(\zeta_{a,L})&x\in(a,L)\\ \zeta_{a,L}(a)=0,\\ \zeta_{a,L}^{\prime}(L)=0,\\ 0\leq\zeta_{a,L}<1\end{cases}$ (27)
Observe that by Lemmas 2.1 and 2.2 and Corollary 2.3, the following holds:
Proposition 4.1.
There exists $b_{0}$ sufficiently large such that if $L-a\geq b_{0}$, then (27) admits a maximum positive solution $\zeta_{a,L}$ such that $0<\zeta_{a,L}<1$ for all $a<x<L$. Moreover, the following are true:
  • For any $\eta>0$ there exists $b=b(\eta)$ so that if La>bL-a>b, then $\zeta_{a,L}(L)>1-\eta$.
  • $\zeta_{a,L}(x)$ is increasing in $x$.
  • If $a+b_{0}\leq L_{1}\leq L_{2}$, then $\zeta_{a,L_{1}}\leq\zeta_{a,L_{2}}$ on $a,L_{1}]$.
  • If $0\leq a_{1}\leq a_{2}\leq L-b_{0}$, then $\zeta_{a_{1},L}\geq\zeta_{a_{2},L}$ on $a_{2},L]$.
The one-dimensional nature of $\zeta_{a,L}$ permits additional analysis. For example, multiplying (27) by $\zeta_{a,L}^{\prime}$ and integrating, we find a quantity that is conserved for all $x$ for a single solution $\zeta_{a,L}$:
$E(a,L)=\frac{d_{u}}{2}(\zeta_{a,L}^{\prime})^{2}+F(\zeta_{a,L}),$ (28)
where $F$ is the antiderivative of $f$:
$F(s)=\int_{0}^{s}f(s^{\prime})\,ds^{\prime}.$ (29)
Several properties of $F$ are important: $F(0)=0$, and due to property F4, there is an additional root $\theta^{\prime}\in(\theta,1)$. For $s\in(0,\theta^{\prime})$, $F(s)<0$, reaching a minimum at $s=\theta$. For $s\in(\theta^{\prime},1)$, $F(s)>0$, reaching a maximum at $s=1$. Thus, for all $y\in(0,F(1)]$, there is a unique solution to $F(s)=y$, and that solution is in $(\theta^{\prime},1)$.
The following analogue of Theorem 1.2 guarantees existence of nontrivial solutions to (26).
@@ 4 Thin-limit analysis in 1D
Consider the system (26) with $f$ of bistable type satisfying the assumptions F1-F4. Assume that $\gamma<\alpha$. There exists $b_{0}$ (at least as large as $b_{0}$ in Proposition 4.1) such that for any $a,L$ such that $L-a\geq b_{0}$, there exists a solution $(u,v)$ of system (26) such that $\zeta_{a,L}<u<1$ in $(a,L)$ and $v>0$ in $(0,a)$. In particular, $u$ is nonconstant.
@@
[…]
Alternatively, Theorem 4.2 could be proved using a reflection argument.
The focus of this section is the behavior of solutions to (26) in the limit as $a\to 0$. One may expect that a predator domain of size zero cannot support any predators, but this is not necessarily the case. Consider, for example, fishermen fishing along the shore of a lake—the shoreline has zero area and the fishermen can still catch fish that come close to the shore. Indeed, our first result on this subject is that as $a\to 0$, the total predator population in solutions to (26) does not vanish and that the prey population density approaches a solution to
$\begin{cases}-d_{u}u^{\prime\prime}=f(u)&0<x<L\\ u(0)=\frac{\gamma}{\alpha}\\ u^{\prime}(L)=0.\end{cases}$ (30)
The system (30) may have many solutions. Indeed, by taking $L$ arbitrarily large, there are arbitrarily many oscillatory solutions with $0<u<\theta^{\prime}$. But the following Proposition 4.3 identifies a unique solution such that $u(L)$ is close to 11 for $L$ large enough, and Theorem 4.4 identifies this as the solution of (30) to which solutions of (26) converge as $a\to 0$. These results and the remainder of the results in this section will require us to add two additional conditions for the function $f:[0,1]\to\mathbb{R}$ to the conditions F1-F4 in Section 1.2:
  • (F5)
    $f$ is continuously differentiable
  • (F6)
    $f^{\prime}(1)<0$.
These conditions will allow us to consider the linearization of $f$ and control the behavior of solutions when $u\approx 1$.
Proposition 4.3.
Suppose $f$ satisfies F1-F6. There exists $\eta_{0}\in(0,1-\theta^{\prime})$ and $b_{0}>0$ (at least as large as $b_{0}$ as in Proposition 4.1) so that if $L>b_{0}$, then there exists a unique solution $u$ to (30) such that $0<u<1$ and $u(L)>1-\eta_{0}$. Moreover, this solution $u$ is monotone increasing and $u(L)$ is increasing in $L$, and approaches 11 as $L\to\infty$.
Equations of the form $d_{u}u^{\prime\prime}=f(u)$ are well studied for a rich variety of functions $f$ (not necessarily satisfying F1-F6). Many results for such equations are found in [23, 46] and references therein. The proof of Proposition 4.3 is similar to that of earlier results (differing mainly in the mixed boundary conditions of (30)), so it is relegated to Appendix A.
Theorem $4.$
Suppose $\alpha>\gamma$. Fix $a_{0}>0$, let $b_{0},\eta_{0}$ be as in Proposition 4.3 and let $L_{0}>a_{0}+b_{0}$ be sufficiently large that $\zeta_{a_{0},L_{0}}(L_{0})>1-\eta_{0}$. For any $0\leq a<a_{0}$ and fixed $L>L_{0}$, let $u_{a}$ and $v_{a}$ denote a coexistence equilibrium solution to (26) guaranteed by Theorem 4.2. Then the following hold in the limit $a\to 0$:
  • (i)
    $u_{a}$ converges in $C^{0}$ to the unique solution $u$ to (30) guaranteed by Proposition 4.3.
  • (ii)
    $u_{a}$ converges to $u$ in $C^{2}$ in every interval $(x_{0},L]$ for $0<x_{0}<L$.
  • (iii)
    $v_{a}$ is unbounded in $C^{0}$, but
    $0<\lim_{a\to 0}\int_{0}^{a}v_{a}(x)\,dx=\frac{d_{u}\alpha}{\beta\gamma}u^{\prime}(0)<\infty.$ (31)
    That is, $v_{a}$ converges to a Dirac mass concentrated at $x=0$.
@@
First, observe that
$\begin{split}\int_{0}^{a}v_{a}\,dx&=\frac{\alpha}{\gamma}\int_{0}^{a}u_{a}v_{a}\,dx\\ &=\frac{\alpha}{\beta\gamma}\int_{0}^{a}\left(d_{u}u_{a}^{\prime\prime}+f(u_{a})\right)\,dx\\ &=\frac{\alpha}{\beta\gamma}\left(-\int_{a}^{L}d_{u}u_{a}^{\prime\prime}\,dx+\int_{0}^{a}f(u_{a})\,dx\right)\\ &=\frac{\alpha}{\beta\gamma}\left(\int_{0}^{L}f(u_{a})\,dx\right)\\ &\leq\frac{\alpha L}{\beta\gamma}\|f\|_{C^{0}}.\end{split}$ (32)
Thus, $v_{a}$ is bounded in $L^{1}$ in the limit $a\to 0$.
Now we show that $u_{a}$ is bounded in $C^{1}$. We know $\zeta_{a,L}<u_{a}<1$, so $u_{a}$ is bounded in $C^{0}$. For $x\in(0,L)$, we calculate
$|u_{a}^{\prime}(x)|$ $\leq\int_{0}^{x}\left|-f(u_{a})+\beta 1_{(0,a)}u_{a}v_{a}\right|\,dx$
$\leq L\|f\|_{C^{0}}+\beta\|v_{a}\|_{L^{1}}.$
Since $v_{a}$ is bounded in $L^{1}$, $u_{a}$ is bounded in $C^{1}$. Thus, by the Arzelà–Ascoli Theorem, the set $\{u_{a}\}$ is precompact in $C^{0}$.
Suppose that $u_{a}$ does not converge to the unique solution $u$ to (30) with $u(L)>1-\eta_{0}$ guaranteed by Proposition 4.3. Then due to the precompactness of $\{u_{a}\}$, there exists a sequence $a_{n}$ in $(0,a_{0})$ and a corresponding sequence $u_{n}:=u_{a_{n}}$ converging to some $\hat{u}\neq u$ in $C^{0}$. For any $x_{0}\in(0,L)$, we have that for $n$ sufficiently large, $a_{n}<x_{0}$ so $d_{u}u_{n}^{\prime\prime}=f(u_{n})$ on $(x_{0},L)$. Integrating twice, we may write for any $x\in(x_{0},L)$,
$u_{n}(x)=u_{n}(x_{0})+u_{n}^{\prime}(x_{0})(x-x_{0})-\frac{1}{d_{u}}\int_{x_{0}}^{x}(x-s)f(u_{n}(s))\,ds.$ (33)
Since $f$ is uniformly continuous and $u_{n}\to\hat{u}$ in $C^{0}$, we conclude that $u_{n}^{\prime}(x_{0})$ converges to some value, say, $d\in\mathbb{R}$. Passing (33) to a limit, we have
$\hat{u}(x)=\hat{u}(x_{0})+d(x-x_{0})-\frac{1}{d_{u}}\int_{x_{0}}^{x}(x-s)f(\hat{u}(s))\,ds.$ (34)
We conclude that $\hat{u}$ is twice differentiable and satisfies